OpenStudy (anonymous):

a square has an area of 49 square inches. if the same amount is added to the length and removed from the width, the resulting rectangle has an area of 45 sgaure inches. find the dimensions of the rectangle. 4in x 9in 3in x 4in 5in x 10in 5in x 9in

OpenStudy (anonymous):

Hi there. Do you just want the answer or learn the material?

OpenStudy (anonymous):

answer please :)

OpenStudy (anonymous):

5x9

OpenStudy (anonymous):

ty hun

OpenStudy (anonymous):

You can solve this in wolfram alpha pretty easily

OpenStudy (anonymous):

how so? i know wolfram decently well

OpenStudy (anonymous):

can i just type the question as i did here?

OpenStudy (anonymous):

what was that?

OpenStudy (anonymous):

that's how i solved your problem

OpenStudy (anonymous):

the original dimension is X, so you know X*X=49. You have a new rectangle, of dimensions A*B, and the area of that is 49. You know A is X minus some length (Y). You know B is X plus that same length (Y). Just put all those facts into wolfram alpha together and the answer pops out, a=5, b=9. THat's how I solved your question.

OpenStudy (anonymous):

thank ya!

OpenStudy (anonymous):

If the area of a square is 49, then the each side length is 7. Let x be the amount to add and subtract from 7. Solve the following for x: (7 + x) (7 - x) = 45 x = 2 9 * 5 = 45 The answer is not in the problem selection list.